OpenStudy (anonymous):

difference quotient for radical 7x+5

OpenStudy (anonymous):

\[f(x) = \sqrt{7x + 5}\] then \[f(x+h) = \sqrt{7(x+h) +5} = \sqrt{7x +7h + 5}\] now subtract f(x) from f(x+h) to get \[f(x+h) - f(x) = \sqrt{7x + 7h + 5} - \sqrt{7x + 5}\] now divide both sides by h to get \[[f(x+h)-f(x)]/h = [\sqrt{7x+7h+5} - \sqrt{7x+5}]/h\] from here, it's going to get a little messy because you will need to multiply the numerator and denominator by \[[\sqrt{7x+7h+5} + \sqrt{7x+5}\] this is called the conjugate. you will have to do some simplification, then apply h approaching zero to get \[[f(x+h)-f(x)]/h = 1/(2\sqrt{7x+5})\] let me know if you need further assistance.

OpenStudy (anonymous):

i appreciate all the work, but i think your final solution is wrong

OpenStudy (anonymous):

i forgot the 7...the answer should be \[7/(2\sqrt{7x+5})\]

OpenStudy (anonymous):

that is still incorrect

OpenStudy (anonymous):

let me put it another way...\[(7/2)(\sqrt{7x+5})^{-1/2}\]

OpenStudy (anonymous):

if it is still wrong...please, rewrite the equation using equation editor

OpenStudy (anonymous):

difference equation of f(x)=\[\sqrt{7x+5}\]

OpenStudy (anonymous):

what solution are you getting...

OpenStudy (anonymous):

\[\sqrt{7h}/5\]

OpenStudy (anonymous):

have you covered derivatives?

OpenStudy (anonymous):

yes, is difference quotient a derivative cuz then i could bypass all this stupid (x+h) stuff

OpenStudy (anonymous):

what i did is basically took the derivative of f(x)...that is the result. difference quotient will eventually turn out to be derivative later on...most likely the next section or chapter.

OpenStudy (anonymous):

ok, i will just try to do it as a derivative then thank you

OpenStudy (anonymous):

yw